# Pruning the Unsurprising: Efficient Code Reasoning via First-Token Surprisal

### v1

@@ Abstract
Recently, Large Reasoning Models (LRMs) have demonstrated remarkable capabilities in code reasoning by scaling up the length of Chain-of-Thought (CoT). However, excessively long reasoning traces introduce substantial challenges in terms of training cost, inference latency, and deployment feasibility. While various CoT compression approaches have emerged to address this challenge, they face inherent trade-offs: token-level methods often disrupt syntactic and logical coherence, while step-level methods based on perplexity fail to reliably capture the logically critical reasoning steps. In this paper, we propose ASAP (Anchor-guided, SurprisAl-based Pruning), a novel coarse-to-fine framework for CoT compression. ASAP first performs anchor-guided pruning to preserve the core reasoning structure, which efficiently reduces the search space for subsequent processing. It then enables a logic-aware pruning by selecting logically essential reasoning steps based on a novel first-token surprisal metric. Finally, ASAP teaches models to autonomously generate and leverage these concise CoTs at inference time, enabling efficient reasoning in coding tasks. Experiments show that ASAP achieves state-of-the-art accuracy across multiple code generation benchmarks while substantially reducing training and inference costs. On the challenging LiveCodeBench v4_v5 benchmark, our approach reduces token generation by 23.5% and inference latency by 43.5% compared to the strongest baseline, while achieving a competitive accuracy of 36.19% in Pass@1. Our results highlight a promising direction for building powerful and efficient LRMs. [^1] Code and model available at https://github.com/Zengwh02/ASAP.
Figure 1: Illustration of CoT pruning by ASAP. The original CoT emitted by LRMs contains redundant branches (highlighted in red dashed boxes). ASAP prunes them in the first stage. It then identifies the first tokens of reasoning steps (marked in blue), whose surprisal scores are computed to guide the second-stage pruning.
@@ 1 Introduction
Recently, the emergence of Large Reasoning Models (LRMs), including OpenAI’s o1 (Jaech et al. 2024) and DeepSeek-R1 (Guo et al. 2025), has demonstrated remarkable capabilities across a wide range of domains. LRMs empower LLMs with the Chain-of-Thought (CoT) prompting paradigm (Wei et al. 2022), which prompts models to “think step by step”. By generating longer reasoning traces during inference, LRMs are capable of handling complex code reasoning tasks (Shi et al. 2024).
Despite achieving significant performance gains, the current LRMs come at a heavy computational cost when the CoTs scale up. Long CoTs often contain substantial redundancy, particularly pronounced in code generation tasks, where LRMs frequently explore diverse strategies or engage in self-debugging (Bi et al. 2024; Xie et al. 2025; Wu et al. 2025; Qu et al. 2025). This not only increases inference and training costs but also suggests a counterintuitive insight: removing redundant reasoning traces may improve both efficiency and performance.
To improve the efficiency of CoT reasoning, a growing body of research has emerged on efficient reasoning (Qu et al. 2025). A prominent line of work is CoT compression for efficient model fine-tuning, which aims to distill concise yet effective reasoning patterns into the model. For example, TokenSkip (Xia et al. 2025a) adapts general-purpose context compressors such as Selective Context (Li et al. 2023) and LLMLingua-2 (Pan et al. 2024). However, these token-level compression methods are fundamentally ill-suited for code reasoning as they damage the logic of an entire code block within the trace, even if only a variable name is removed. The compressed context makes it hard for models to learn effective reasoning patterns. A more promising direction is step-level pruning. For example, SPIRIT (Cui et al. 2025) prunes the reasoning steps while preserving structural and semantic integrity. However, these approaches face a deeper challenge: How to reliably estimate the logical importance of a reasoning step. SPIRIT relies on perplexity (PPL), but we argue that PPL is an inadequate metric as it primarily reflects linguistic fluency rather than logical necessity and has shown limitations in long-context scenarios (Hu et al. 2024; Fang et al. 2024).
To overcome these challenges, we propose ASAP: A novel framework for Anchor-guided, SurprisAl-based Pruning of CoTs. Our core insight, illustrated in Figure 1, is identifying and pruning redundant steps in CoTs while preserving a coherent logical path. In the first stage, ASAP leverages the inherent structural and logical nature of code to generate a step-by-step reasoning path. The path serves as an anchor for coarse-grained pruning, efficiently identifying and preserving the core reasoning backbone. In the second stage, it enables a fine-grained logic-aware pruning by selecting the first tokens of reasoning steps based on a new First-Token Surprisal metric. The resulting CoT retains only the steps that introduce novel and meaningful information. By fine-tuning on the pruned CoTs, we teach a large reasoning model to generate and leverage these compact reasoning patterns at inference time, enabling efficient reasoning in coding tasks.
@@
We validate our approach through extensive experiments on the DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B (Guo et al. 2025) across a suite of challenging code generation benchmarks, including HumanEval (Chen et al. 2021), HumanEval+ (Liu et al. 2023), LiveCodeBench (Jain et al. 2024), and LeetCodeDataset (Xia et al. 2025b). Our resulting model establishes a new state-of-the-art in the performance-efficiency trade-off. Notably, on the challenging LiveCodeBench v4_v5 benchmark, ASAP achieves 36.19% Pass@1 while reducing token generation by 23.5% and inference latency by 43.5% compared to the strongest baseline.
@@
  • We propose ASAP, a novel two-stage pruning framework that combines coarse-grained anchor-guided pruning with fine-grained surprisal-based pruning to identify and retain logically critical steps.
  • We are the first to propose and empirically validate First-Token Surprisal as a superior metric to perplexity for measuring the logical importance of CoT steps in the code generation task.
@@
  • Extensive experiments on multiple code generation benchmarks demonstrate that models fine-tuned on CoTs pruned by ASAP achieve state-of-the-art accuracy while substantially reducing both training and inference costs.
2 Related Work
2.1 Chain-of-Thought and Advanced Reasoning
Chain-of-Thought (CoT) prompting has emerged as a foundational technique for eliciting reasoning in large language models (LLMs) (Wei et al. 2022). This has sparked a surge of research exploring more sophisticated reasoning frameworks such as Tree of Thoughts (Yao et al. 2023), Graph of Thought (Lei et al. 2023), Natural Program (Ling et al. 2023), etc. More recently, the emergence of powerful Large Reasoning Models (LRMs) such as OpenAI’s o1 (Jaech et al. 2024) and DeepSeek-R1 (Guo et al. 2025) has further advanced the field. These models exhibit strong long-chain reasoning capabilities, which are often enhanced through supervised fine-tuning (SFT) and reinforcement learning (RL) (Jaech et al. 2024; Guo et al. 2025; Kimi et al. 2025; Yang et al. 2025; Yu et al. 2025; Wang et al. 2025), especially in domains such as mathematical and programming reasoning. Unlike prior work that focuses on enhancing reasoning by generating longer or more complex CoTs, we aim to improve both reasoning performance and efficiency by pruning redundant steps from existing CoTs.
@@ 2.1 Chain-of-Thought and Advanced Reasoning
Figure 2: The overall framework of ASAP. The process begins with a training sample consisting of a question, an intermediate CoT, and the answer. In Stage 1, an LLM generates a “Direct CoT” ($C_{direct}$) from the question and answer. $C_{direct}$ acts as an anchor to prune the “Original CoT” ($C_{origin}$) into a “Coarse-grained Pruned CoT” ($C_{coarse}$). In Stage 2, we compute the first-token surprisal for each step in $C_{coarse}$ to remove less critical steps, yielding the final “Fine-grained Pruned CoT” ($C^{\prime}$) for fine-tuning.
2.2 Context Compression for LLMs
As tasks become more complex, the context in LLMs has expanded dramatically, resulting in significant computational overhead. This has motivated a line of work on context compression (Li et al. 2025; Fang et al. 2025), aiming to reduce input length while preserving key information. For instance, Selective Context (Li et al. 2023) leverages self-information to delete redundant lexical units. LLMLingua (Jiang et al. 2023) employs an LLM to select key demonstrations and then applies token-level filtering across the context. Its extension, LongLLMLingua (Jiang et al. 2024), introduces document reordering and subsequence reconstruction to enable compression over longer contexts. LLMLingua-2 (Pan et al. 2024) further advances this line by distilling training data and training a token classifier for task-agnostic compression. While these methods have proven effective for text compression, they are not tailored to the structured and logically coherent nature of CoT reasoning. In contrast, our method is explicitly designed for CoT compression, which identifies and preserves logically essential reasoning steps.
2.3 Efficient Reasoning via Fine-Tuning
The increasingly lengthy reasoning traces impose substantial costs, which have motivated a growing body of work focused on efficient reasoning (Qu et al. 2025). Several recent approaches aim to compress CoTs while preserving their effectiveness. C3oT (Kang et al. 2025) uses GPT-4 as a high-quality external compressor to generate shorter CoTs that retain essential reasoning content. TokenSkip (Xia et al. 2025a) enables controllable compression by identifying and removing less important tokens. SPIRIT (Cui et al. 2025) identifies important reasoning steps by detecting perplexity shifts when individual steps are removed. Another paradigm involves compressing CoTs into continuous latent representations, as explored by Coconut (Hao et al. 2024), CCoT (Cheng and Van Durme 2024), CODI (Shen et al. 2025), etc. Our method, ASAP, distinguishes itself from these methods through its explicit two-stage pruning framework, which first generates an anchor for coarse-grained pruning, followed by a fine-grained refinement based on first-token surprisal. By explicitly pruning at the step level, ASAP produces logically condensed human-readable CoTs.
@@ 3.1 Overall Framework
Given a training corpus of triples $\{(Q_{i},\mathbf{C}_{i},A_{i})\}_{i=1}^{N}$—where each $Q_{i}$ denotes a question, $\mathbf{C}_{i}=\{s_{1},\dots,s_{N}\}$ is the intermediate reasoning steps, known as chain-of-thoughts (CoT), generated by a large reasoning model $\mathcal{L}$, and $A_{i}$ represents the predicted answer to the question. Our goal is to compress each $\mathbf{C}_{i}$ into a concise one $\mathbf{C}_{i}^{\prime}$ such that $|\mathbf{C}_{i}^{\prime}|\ll|\mathbf{C}_{i}|$ while the LRM maintains the quality of generated reasoning steps and answers when fine-tuned on the compressed triplets $\{Q_{i},\mathbf{C}_{i}^{\prime},A_{i}\}_{i=1}^{N}$.
We propose ASAP, a novel method for CoT compression, as illustrated in Figure 2. ASAP adopts a coarse-to-fine pruning framework that firstly removes redundant steps from a given reasoning trace and then refines it by selecting the most critical logical steps. In the first stage called Anchor-guided Pruning, ASAP infers a concise, anchor reasoning path called “Direct Thoughts” ($C_{direct}$) based on the $(Q,A)$ pairs. Guided by $C_{direct}$ as an anchor, ASAP prunes the original CoT $C$ into a “coarse-grained pruned CoT” ($C_{coarse}$) using an LLM, which retains the main logical backbone while preliminarily removing digressions and redundancies. In the second stage called Surprisal-based Refining, ASAP performs a logic-aware refinement to $C_{coarse}$. A new metric called First-Token Surprisal is designed to measure the logical importance of a reasoning step and then iteratively filter steps with low surprisal, yielding the final CoT (denoted as $\mathbf{C}^{\prime}$ in our formulation). Finally, all $Q,C^{\prime},A$} tuplets are used to fine-tune a large reasoning model for downstream tasks, which internalizes concise reasoning patterns and substantially improves inference efficiency.
@@ 3.2 Anchor-guided Pruning
Generating the Direct CoT Anchor
We prompt an LLM to generate a concise and coherent anchor reasoning trace, $C_{direct}$, based on the question and answer. An anchor is defined as a structured, step-by-step explanation that outlines how to derive the answer from the question. This leverages the inherent logical and procedural nature of code reasoning. The resulting $C_{direct}$ serves as a logical backbone—an anchor—to guide the subsequent pruning of the more elaborate $C_{origin}$. The prompt used for the generation is shown in Appendix E.
Pruning with Pattern Matching
@@
Using $C_{direct}$ as a reference, we prompt the LLM to prune the original CoT $C$. Specifically, the LLM is instructed to: 1) remove unnecessary reasoning steps from $C$; 2) retain all key supporting content that aligns with the logic of $C_{direct}$; and 3) crucially, preserve the original wording without introducing new information. The prompt used for pruning is shown in Appendix E. To mitigate LLM hallucination, we incorporate a validation step based on Gestalt Pattern Matching (Black 2004), which validates structural and semantic alignment with $C$.

### v2

# Pruning the Unsurprising: Efficient LLM Reasoning via First-Token Surprisal

@@ Abstract
Large Reasoning Models (LRMs) have demonstrated remarkable capabilities by scaling up the length of Chain-of-Thought (CoT). However, excessively long reasoning traces pose substantial challenges for training cost and inference latency. While various CoT compression approaches have emerged to address this challenge, they face inherent trade-offs: token-level methods often disrupt syntactic and logical coherence, while step-level methods based on perplexity fail to reliably capture the logically critical reasoning steps because of the dilution of logical information. In this paper, we propose ASAP (Anchor-guided, SurprisAl-based Pruning), a novel coarse-to-fine framework for CoT compression. ASAP first performs anchor-guided pruning to preserve the core reasoning structure, which efficiently reduces the search space for subsequent processing. Leveraging the insight that logical branching choices are concentrated at the onset of reasoning steps, it then enables logic-aware pruning by selecting logically essential reasoning steps based on a novel first-token surprisal metric. Finally, ASAP distills the models to autonomously generate and leverage these concise CoTs at inference time, enabling efficient reasoning. Experiments show that ASAP achieves state-of-the-art accuracy across multiple benchmarks while substantially reducing training and inference costs.
Pruning the Unsurprising: Efficient LLM Reasoning via First-Token Surprisal
Wenhao Zeng1   Yaoning Wang2   Chao Hu1   Yuling Shi1 Chengcheng Wan3   Hongyu Zhang4   Xiaodong Gu1 1Shanghai Jiao Tong University  2Fudan University 3East China Normal University  4Chongqing University {zengwh_cs, xiaodong.gu}@sjtu.edu.cn
Figure 1: Illustration of CoT pruning by ASAP. The Original CoT generated by LRMs exhibits two types of redundancy: (1) Structural Redundancy, such as digressive branches (highlighted in red dashed boxes), which are removed by our Stage 1 Anchor-guided pruning; and (2) Logical Redundancy within valid paths. ASAP addresses the latter in Stage 2 by computing the surprisal of the first tokens of reasoning steps (marked in blue) to identify and retain only the critical cognitive pivots.
@@ 1 Introduction
The emergence of Large Reasoning Models, including OpenAI’s o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025), marks a paradigm shift in artificial intelligence. By scaling up Chain-of-Thought (CoT) reasoning Wei et al. (2022), these models demonstrate emergent capabilities in complex domains such as mathematics Sun et al. (2025), programming Shi et al. (2024); Yang et al. (2025b); Hu et al. (2025), and logical reasoning Liu et al. (2025); Zhang et al. (2025c, b). However, this performance comes at a prohibitive cost: reasoning traces often span thousands of tokens, introducing substantial latency and memory overhead. Crucially, these lengthy traces often contain substantial redundancy, such as over-explaining simple problems or superficially exploring multiple paths for complex ones Bi et al. (2024); Xie et al. (2025); Wu et al. (2025); Qu et al. (2025). For instance, the Original CoT in Figure 1 contains tangential branches (highlighted in red dashed boxes), such as exploring an alternative manual implementation that is subsequently rejected (“But that’s more code…”). Furthermore, the reasoning is punctuated by syntactic fillers that contribute little to the core logic. This observation raises a fundamental question: Can we identify and retain only the “cognitive pivots” of reasoning while discarding the redundancy?
A growing body of research has emerged on CoT compression for efficient reasoning Qu et al. (2025). Token-level methods like TokenSkip Xia et al. (2025a) adapt general-purpose context compressors such as LLMLingua-2 Pan et al. (2024) to prune non-informative tokens. However, indiscriminate token removal risks disrupting the syntactic integrity of the reasoning chain. To address this, step-level pruning methods like SPIRIT Cui et al. (2025) trim entire reasoning steps, thereby preserving structural coherence. Nevertheless, these approaches face a fundamental challenge: accurately estimating the logical importance of each step. They typically rely on fixed metrics like perplexity (PPL), which measures the overall predictability of a sentence. This holistic measure often dilutes the signal of critical logical leaps with the noise of syntactically predictable but logically trivial content.
In this work, we ground CoT compression from an information-theoretic perspective. Through an empirical analysis of 10 million reasoning tokens (detailed in Section 2), we find that the logical progression within a CoT sequence is not uniformly distributed; instead, its information density is highly concentrated at the beginning of each reasoning step—specifically within the first few tokens (blue-highlighted in Figure 1). These tokens, such as “But” (self-correction) or “So” (deduction, not continuation), serve as high-entropy cognitive pivots. By leveraging the surprisal of these initial tokens, we can distinguish between critical logical transitions and predictable elaborations.
Guided by this insight, we propose ASAP (Anchor-guided, SurprisAl-based Pruning), a coarse-to-fine framework designed to preserve these high-information steps. ASAP in a two-stage cascade that directly addresses the two types of redundancies identified in our study (illustrated in Figure 1): First, it employs Anchor-guided Pruning to remove structural redundancies. By generating a concise step-by-step reasoning trace as a logical backbone, it identifies and prunes the irrelevant branches (e.g., the red boxes in Figure 1). Second, it performs Surprisal-based Refining to eliminate logic-sparse steps. Leveraging our First-Token Surprisal metric, this stage iteratively filters out steps acting as mere fillers while retaining the high-surprisal cognitive pivots. Finally, we distill these compact, logic-dense CoTs into a target model, enabling it to generate efficient reasoning chains.
@@
We validate our approach through extensive experiments on the DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B Guo et al. (2025) across diverse domains. The results demonstrate that ASAP establishes a superior Pareto frontier between performance and efficiency. Notably, on the challenging LiveCodeBench v4_v5 benchmark, ASAP achieves 36.19% Pass@1 while reducing token generation by 23.5% and inference latency by 43.5% compared to the strongest baseline.
@@
  • We present an empirical analysis of the information concentration of CoTs, uncovering that the surprisal of the starting token for each CoT step is a more robust indicator of logical importance than perplexity.
  • We propose ASAP, a novel CoT compression framework that combines structural alignment with information-guided refinement.
@@
  • Extensive experiments on multiple benchmarks demonstrate that models fine-tuned on CoTs pruned by ASAP achieve state-of-the-art accuracy while substantially reducing computational costs.
(a) Entropy Distribution
(b) High-Frequency First Tokens
(c) High-Entropy Tokens
Figure 2: Empirical analysis of 10M tokens from DeepSeek-R1-Distill-Qwen-32B. (a) The entropy distribution reveals a clear information concentration: first tokens (blue) exhibit significantly higher uncertainty (entropy) compared to body tokens (orange), which are highly deterministic. (b) The most frequent first tokens are a mixture of logical operators (e.g., Wait) and ubiquitous syntactic connectors (e.g., So). (c) High-entropy states filter out predictable fillers like So or Then, while exclusively highlighting cognitive pivots such as Perhaps, What, and Alternative.
2 Empirical Analysis
To investigate the intrinsic distribution of logical information in CoTs, we conducted a large-scale analysis on 10 million tokens generated by DeepSeek-R1-Distill-Qwen-32B Guo et al. (2025) across diverse reasoning benchmarks (AIME and LiveCodeBench Jain et al. (2024)).
Information Concentration in CoTs.
We analyze the entropy distribution of the first token of each reasoning step compared to all subsequent tokens. Entropy, in this context, quantifies the model’s uncertainty regarding the next state transition Shannon (1948); Malinin and Gales (2020); Kuhn et al. (2023); Wang et al. (2025); Cheng et al. (2025). As illustrated in Figure 2(a), a distinct concentration is observed. Starting tokens (blue) exhibit a dispersed distribution with a significantly higher 90th percentile. In contrast, other tokens (orange) are heavily concentrated near zero entropy, indicating that once a reasoning step is initiated, its subsequent elaboration is largely deterministic and syntactically driven. This empirical evidence confirms that the logical branching points, where the model actively deliberates on the reasoning path, are structurally concentrated at the beginning of each step.
Identifying Cognitive Pivots with Entropy.
Having identified the informative start tokens, we perform a more in-depth analysis of real logical pivots among the start tokens. We aim to distinguish between superficial syntactic connectors and real logical pivots. Figure 2(b) presents high-frequency start tokens that appear in CoTs, which mix connectors (“So”, “Let”) with reasoning markers (“Wait”, “But”). However, when focusing on tokens generated in high-entropy states (Figure 2(c)), a qualitative shift emerges. Predictable connectors like “So” and “Then” are effectively suppressed due to their low uncertainty. Instead, the distribution is dominated by terms representing cognitive pivots and state transitions, such as: 1) Exploration: “Alternative”, “Another” (proposing hypotheses or new paths). 2) Causality: “Because”, “Since” (providing formal justification). 3) Self-Correction: “Perhaps”, “What” (indicating error detection or logic reversal).
This analysis demonstrates that high entropy is a robust indicator of logical salience. Since the actual next token is known in the given training sequence, we operationalize this insight by using First-Token Surprisal as a proxy to identify and preserve these critical reasoning hops in our proposed framework Fu et al. (2025).
@@ 2 Empirical Analysis
Figure 3: The overall framework of ASAP. The pipeline consists of three phases: (1) In Stage 1, the LLM generates a “Direct Thought" ($\mathcal{P}$) from the (Question, Answer) pair. $\mathcal{P}$ acts as an anchor to prune the “Original CoT" ($C$) into a “Coarse-grained Pruned CoT" ($C_{coarse}$). (2) In Stage 2, we compute the First-Token Surprisal for each step in $C_{coarse}$. High-surprisal steps are retained, while low-surprisal fillers are pruned, yielding the final “Fine-grained Pruned CoT" ($C^{\prime}$). (3) In Training Stage, the data with ASAP pruned CoTs is used to fine-tune the LRM for efficient inference.
@@ 3.1 Overall Framework
Formally, we consider a supervised reasoning task defined by a dataset $\mathcal{D}=\{(Q_{i},C_{i},A_{i})\}_{i=1}^{N}$, where $Q_{i}$ is the query, $A_{i}$ is the predicted answer, and $C_{i}$ represents the original CoT generated by the LRMs. $C_{i}$ is a sequence of reasoning steps $C_{i}=\{s_{1},s_{2},\dots,s_{L}\}$. Our goal is to compress each $C_{i}$ into a concise pruned one $C_{i}^{\prime}$ such that $|C_{i}^{\prime}|\ll|C_{i}|$ while the LRMs maintains the quality of generated reasoning steps and answers when fine-tuned on the dataset $\mathcal{D^{\prime}}=\{(Q_{i},C_{i}^{\prime},A_{i})\}_{i=1}^{N}$.
We propose ASAP, a coarse-to-fine framework tailored to the redundancy of “Original CoT" ($C$), as illustrated in Figure 3. Stage 1 (Anchor-guided Pruning) reduces structural redundancy (e.g., dead ends) by aligning the CoT with a generated logical backbone. The LLM generates a “Direct Thought" ($\mathcal{P}$) from the $(Q,A)$ pairs. $\mathcal{P}$ acts as an anchor to prune the $C$ into a “Coarse-grained Pruned CoT" ($C_{coarse}$). Stage 2 (Surprisal-based Refining) reduces logical redundancy (e.g., syntactic fillers) by filtering non-informative steps. We approximate the information of each step in $C_{coarse}$ with their surprisal of start tokens and prune low-surprisal steps, yielding the final “Fine-grained Pruned CoT" ($C^{\prime}$). Finally, all $\{(Q,C^{\prime},A)\}$ are utilized to fine-tune the target model.
@@ 3.2 Anchor-guided Pruning
Directly pruning raw CoTs is challenging due to the noise and unstructured digressions inherent in LLM reasoning Zhou et al. (2024). To address this, we first construct a high-level logical skeleton to narrow the pruning space.
Generate Direct Thoughts.
We prompt the LLM to infer a concise reasoning path called “Direct Thought” ($\mathcal{P}$) based on the $(Q,A)$ pair (see Appendix B for prompts). Unlike exploratory CoTs, $\mathcal{P}$ is generated as a structured, step-by-step explanation that outlines how to derive the answer from the question, exemplified in Appendix B. This $\mathcal{P}$ acts as a reference anchor, outlining the least reasoning trajectory required to solve the problem.
Pruning with Pattern Matching.
@@
Guided by the anchor $\mathcal{P}$, we prompt the LLM to prune the original CoT $C$. Specifically, the LLM is instructed to: 1) remove unnecessary reasoning steps from $C$; 2) retain all key supporting content that aligns with the logic of $\mathcal{P}$; and 3) crucially, preserve the original wording without introducing new information. The prompt used for pruning is shown in Appendix B. The goal is to extract the subsequence of $C$ that semantically aligns with $\mathcal{P}$ while discarding irrelevant branches (as shown in the “Original CoT” block of Figure 3).